When solutions of AgNO₃ and NaOH react, the balanced molecular equation is: 2 AgNO₃(aq) + 2 NaOH(aq) → Ag₂O(s) + 2 NaNO₃(aq) + H₂O(l) How much Ag₂O is produced when 0.200 g of AgNO₃ and 0.200 g of NaOH react?
1
Expert's answer
2020-12-06T13:54:22-0500

2AgNO₃(aq) + 2NaOH(aq) → Ag₂O(s) + 2NaNO₃(aq) + H₂O(l)


n=m/M

n(AgNO₃)=0.200 g / 170 g/mol = 0.00118 mol

n(NaOH)=0.200 g / 40 g/mol = 0.005 mol

n(AgNO₃)<n(NaOH)"\implies" lack of AgNO₃


n(Ag₂O)= 1/2n(AgNO₃) = 5.9*10-4 mol

m(Ag₂O)=n*M=5.9*10-4 mol * 232 g/mol = 0.137 g